Mg(OH)2 + 2HCl = MgCl2 + 2H2O

Assume that a person takes 0.100 g antacid. How many mol of MgCl will be produce from the ingested antacid, assuming all of the antacid undergoes the reaction?
1
Expert's answer
2020-11-17T10:45:00-0500

Mg(OH)2 is the antacid; m(Mg(OH)2) = 0.100 g;

Molar mass Mr(Mg(OH)2) = Ar(Mg) + 2Ar(O) + 2Ar(H) = 24.305 g/mol + 2*16 g/mol + 2*1 g/mol = 58.305 g/mol.

n(Mg(OH)2) = m(Mg(OH)2)/Mr(Mg(OH)2) = (0.100 g)/(58.305 g/mol) = 0.0017 mol;

n(MgCl2) = n(Mg(OH)2) (stoichiometric coefficients);

n(MgCl2) = 0.00172 mol.